OpenStudy (anonymous):

simplify a^6/a^2 a)-a/a^3 b)1/a^4 c)a^4 d)a^5/a

jimthompson5910 (jim_thompson5910):

Rule: \[\Huge \frac{x^y}{x^z} = x^{y-z}\]

jimthompson5910 (jim_thompson5910):

when you divide two exponential expressions with the same base, you subtract the exponents

OpenStudy (anonymous):

so a^4 is the answer

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

i have a couple more questions

OpenStudy (anonymous):

identify the property demonstrated: (7*3)^4=7^4*3^4 a)power of a power property b)quotient of powers property c)product of powers property d)power of a product property

jimthompson5910 (jim_thompson5910):

what is 7*3? a power? a product?

OpenStudy (anonymous):

is D the answer

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

simplify (w/6)^2 a)w/36 b)w^2/6 c)w^2/36 d)36w^2